If I have eight test. What's the lowest score I can get to average and 89.
The lowest grade is 0 and the highest is a 100

Answer by stanbon(75887)

If I have eight test. What's the lowest score I can get to average and 89.
The lowest grade is 0 and the highest is a 100
----------
Assume you get 100 on 7 of the tests
Find the score of the 8th test.
Average = (7*100+x)/8 = 89
-------
700+x = 712
x = 12
-----
Ans: 12 is the lowest score you can get and still average 89
---------------
cheers,
Stan H.
-------------